OpenStudy (agl202):

Solve... −2x − 13 = 8x + 7. 1.) −2 2.) 2 3.) −6 4.) 6

Mehek (mehek14):

\(-2x-13=8x+7\\8x+2x=?\)

OpenStudy (agl202):

10x

OpenStudy (spring98):

ya

Mehek (mehek14):

correct \(-13=10x+7\\-13-7 = ?\)

OpenStudy (agl202):

-20

Mehek (mehek14):

\(-20=10x\) you would have to divide \(\dfrac{-20}{10}=x\)

OpenStudy (agl202):

x= -2

Mehek (mehek14):

good job!

OpenStudy (agl202):

Thx...

Mehek (mehek14):

yw ^_^